Facebook recently examined all active Facebook users and determined that the average user has 183 Facebook friends. Suppose that the population of number of Facebook friends of a randomly selected user has a standard deviation of \sigmaσ = 241. Facebook draws a SRS of 250 Facebook users. What is the probability that the average number of Facebook friends is less than 175? Apply the Central Limit Theorem.
Answer #1

Given,

= 183 , = 241

Using central limit theorem,

P( < x) = P(Z < x - / )

So,

P( < 175) = P(Z < 175 - 183 / (241 / sqrt(250) ) )

= P(Z < -0.5249)

= 1 - P(Z < 0.5249)

= 1 - 0.7002

= 0.2998